Suppose you borrow $200 from a payday lender for one week at a weekly rate of 10%. You'll obviously owe $220 at the end of a week. If you are unable to repay the loan, however, the lender will say that you now owe not only $220 but also 10% of that $220 at the end of the second week.

How much will you owe by the end of the second week? $

If you still can't repay any of it, how much will you owe at the end of the third week? $
Answer #1

Amount owed at end of 2nd week = 220*(1+10%) = $242

Amount owed at the end of 3rd week = 242* (1+10%) = $266.20